OpenStudy (anonymous):

TRIGONOMETRY: If the central angle theta =10rad, and the area of the sector formed is 25square units, find the AREA of the circle. TRIGONOMETRY: If the central angle theta =10rad, and the area of the sector formed is 25square units, find the AREA of the circle. @Mathematics

OpenStudy (ash2326):

sorry Area of sector =pi*r^2*theta/2pi 25=pi*r^2*10/2pi pi*r^2=25*2pi/10 Area of circle=15.7

OpenStudy (anonymous):

\[\exp\]

OpenStudy (anonymous):

Area of circle=15.7